Find fx(x,y), fy(x,y), fx(1,3), and fy(-2,4) for the given function. If

𝑧 = 𝑓(π‘₯, 𝑦) = 3π‘₯ଷ𝑦ଢ - π‘₯ଢ𝑦ଷ + 4π‘₯ + 9


1
Expert's answer
2021-06-09T15:38:23-0400

βˆ‚x. (x, y) ≑ fx(x, y) ≑ Dxf(x, y) ≑ f1;. β€’ partial derivative of f with respect to y is ... to find fy(x, y): keeping x constant, take y derivative. ... Functions of n > 2 Variables: given f(x) = f(x1,..., xn).

A function of one variables has a single rate of change then, F(x) (a, b)) is the rate of change of F with respect to x ( respect to y) at the simplarly, Fy (1,3) is the derivative of v(y)= f(1,y) at = y= 3(6)-6(0)+4+9=18-0+13=31

function one variables has a single rate of change then, F(x) (a, b)) is the rate of change of F with respect to x ( respect to y) at the simplarly, Fy (1,3) is the derivative of v(y)= f(1,y) at = y= 3(6)-6(0)+4+9=18-0+13=31

If

𝑧 = 𝑓(π‘₯, 𝑦) = 3π‘₯ଷ𝑦ଢ - π‘₯ଢ𝑦ଷ + 4π‘₯ + 9= F(g(3+4+9)====> F(g(16) ans